Suppose researchers interested in the effects of adolescent exposure to mold in a household is associated with asthma. Researchers collected information about a sample of children living in homes determined to have a significant amount of potential mold exposure and determined that 154 out of the 213 children had asthma. After a sample of 428 households with not a sufficient amount of mold exposure, only 108 children were found to have asthma.

A) Conduct a statistical test to determine if there is a significant risk difference for asthma based on exposure to mold at the alpha level of 0.05.

B) Construct a confidence interval about your point estimate for the difference in risk between the groups.

PLEASE CHOOSE THE MOST APPROPRIATE OPTION ( Please do not post calculations that do not match the options given, I must choose one of the options and I cannot get to the right answer, thank you.

Option 1

A) risk difference=0.471

  Ho: p1=p2; Ha: p1≠p2

Z stat=12.7

Pvalue-<0.0001

B) CI (0.40, 0.54)


Option 2

A) risk difference=0.211

Ho: p1=p2; Ha: p1≠p2

Z stat=17.2

Pvalue-<0.05

B)CI (0.23, 0.56)

Option 3

A) risk difference=1.741

  Ho: p1=p2; Ha: p1≠p2

Z stat=10.7

Pvalue-<0.0001

B) CI (0.04, 0.054)
Answer #1
mold no mold
x1                =    154 x2                =    108
p̂1=x1/n1 = 0.7230 p̂2=x2/n2 = 0.2523
n1                       = 213 n2                       = 428
estimated prop. diff =p̂1-p̂2    = 0.471
std error Se=√(p̂1*(1-p̂1)*(1/n1+1/n2) = 0.0412
test stat z=(p̂1-p̂2)/Se = 12.7

option A is correct

Option 1

A) risk difference=0.471

  Ho: p1=p2; Ha: p1≠p2

Z stat=12.7

Pvalue-<0.0001

B) CI (0.40, 0.54)